OpenStudy (anonymous):

30(i1-i2)-15i2=0 30i1+15i2=24

OpenStudy (anonymous):

posted

OpenStudy (lilg132):

equation 1) 30i1 - 30i2 - 15i2 = 0 30i1 - 45i2 = 0 equation 2) 30i1 + 15i2 = 24 30i1 = 24 - 15i2 equation 1) 24 - 15i2 - 45i2 = 0 (I am substituting here) -60i2 = -24 24 = 60i2 24/60 = 60i2/60 0.4 = i2 now we now what is i2 we can work out i1 equation 2) 30i1 + 15(0.4) = 24 30i1 + 6 = 24 30i1 = 24 - 6 30i1 = 18 30i1 / 30 = 18 / 30 i1 = 0.6 Dont forget to click best response !

OpenStudy (anonymous):

thanks so how well do u know matrices